OpenStudy (anonymous):

Rationalize : 6 / root 12 - root 3 ! Pls show wid steps , got stuck in between !

jhonyy9 (jhonyy9):

so sqrt12 = sqrt(4*3) = sqrt(2^2 *3) = 2sqrt3 yes ? so than the denominator will be in this case 2sqrt3 - sqrt3 = sqrt3 do you can continue it now ?

OpenStudy (anonymous):

You itself continue , and finish

jhonyy9 (jhonyy9):

6/sqrt3 = 6sqrt3 /3 = 2sqrt3 do you understand it ?

OpenStudy (anonymous):

Ys , tnx a looot !

jhonyy9 (jhonyy9):

my pleasure